Betty wants to install new carpet in her living room. The room length is 15 ft and the width is 15ft. If the cost of the carpet is $8 per square foot, what will be the total cost to carpet Betty's living room? After installing the carpet, Betty decides to also install floor wood molding for her living room. The molding will go around the entire room. The molding costs $4 per foot. What will be the total cost of the molding?
1
Expert's answer
2019-11-14T09:14:19-0500

Room area: A=15*15=25 sq.ft.

Cost of the carpet: C=A*$8=225*$8=$1800.

Room perimeter: P=(15+15)*2=60 ft.

Cost of the molding: M=P*$4=60*$4=$240.